# Population Transformer: Learning Population-level Representations of Neural Activity

### v2

@@ Abstract
We present a self-supervised framework that learns population-level codes for arbitrary ensembles of neural recordings at scal. We address two key challenges in scaling models with neural time-series data: sparse and variable electrode distribution across subjects and datasets. The Population Transformer (PopT) stacks on top of pretrained representations and enhances downstream decoding by enabling learned aggregation of multiple spatially-sparse data channels. The pretrained PopT lowers the amount of data required for downstream decoding experiments, while increasing accuracy, even on held-out subjects and tasks. Compared to end-to-end methods, this approach is computationally lightweight and more interpretable, while still retaining competitive performance. We further show how our framework is generalizable to multiple time-series embeddings and neural data modalities. Beyond decoding, we interpret the pretrained PopT and fine-tuned models to show how they can be used to extract neuroscience insights from massive amounts of data. We release our code as well as a pretrained PopT to enable off-the-shelf improvements in multi-channel intracranial data decoding and interpretability.
@@ 1 Introduction
Building effective representations of neural data is an important tool in enabling neuroscience research. We are particularly interested in modeling neural activity recorded from spatially varying electrode layouts, such as intracranial recordings, which use invasive probes placed within the brain to provide high temporal resolution recordings of local neural activity (Parvizi & Kastner, 2018; Herff et al., 2020). But because of dispersed electrode placement within the brain volume, intracranial recordings (iEEG) suffer from data sparsity, and there is significant variability in probe placement across subjects (Parvizi & Kastner, 2018; Herff et al., 2020), leading to high variability in the semantics of the input channels. While these difficulties are particularly relevant for iEEG, they also exist for other neural data modalities such as non-invasive EEG due to variation in electrode layouts. Furthermore, annotated training data is often scarce, since it is limited by patient availability and labor-intensive labeling. This is a problem for many existing approaches which rely on supervised learning (Faezi et al., 2021; Herff et al., 2020; Martin et al., 2018; Metzger et al., 2023; Willett et al., 2023).
To improve decoding data-efficiency, self-supervised pretraining on unannotated data can be employed to first learn generic representations of the recordings, independent of the downstream decoding task that requires task labels. This means that the model does not have to use valuable annotated samples to learn how to do feature extraction before it can do classification, improving the reach of neuroscientific research. In this paper, we are interested in developing generic representations of multi-channel intracranial recordings that enable efficient adaptation to a wide range of downstream decoding tasks. Prior work has shown how to pretrain subject-specific (Le & Shlizerman, 2022) or channel-specific (Wang et al., 2022) models of intracranial data, but such techniques ignore inter-channel relationships or commonalities that might exist across subjects. The most general approach would be to pretrain using data from multiple datasets, but would require tackling the aforementioned challenges of sparse electrode coverage and variable electrode placement between subjects.
@@
We propose Population Transformer (PopT), a self-supervised pretraining approach that learns subject-generic representations of arbitrary electrode ensembles. Transformers offer the flexibility to learn aggregate information across channel configurations, but massive data is needed to train the attention weights (Devlin et al., 2019). During pretraining, we train on large amounts of unannotated data and simultaneously optimize both a channel-level and ensemble-level objective. This requires the model to (1) build subject-generic representations of channel ensembles and (2) meaningfully distinguish temporal relationships between different ensembles of channels.
@@
Our PopT approach is modular, and builds on top of powerful single-channel temporal embeddings, which provides two key advantages. First, by separating the single-channel embedding and multi-channel-aggregation into different modules, we make our approach agnostic to the specific type of temporal embedding used, leaving room for future independent improvements along either the temporal or spatial dimension (an approach that has been validated in video modeling (Arnab et al., 2021)). Second, by taking advantage of learned channel embeddings, PopT training is computationally lightweight compared to their end-to-end counterparts (Appendix B) and baseline aggregation approaches (Figure 4), allowing for adoption in lower compute resource environments.
@@
Empirically, we find that our pretrained PopT outperforms commonly used aggregation approaches (Ghosal & Abbasi-Asl, 2021), and is competitive with end-to-end trained methods (Zhang et al., 2024; Yang et al., 2024; You et al., 2019). Moreover, we find that these benefits hold even for subjects not seen during pretraining, indicating its usefulness for new subject decoding. We also show that the pretrained PopT weights themselves reveal interpretable patterns for neuroscientific study. Finally, we demonstrate that our proposed framework is agnostic to the underlying temporal encoder further allowing it to adapt to other neural recording modalities.
@@ 2 Related Work
Modeling across variable input channels Modeling spatial representations on top of temporal embeddings has been found to be beneficial for decoding (Faezi et al., 2021; Le & Shlizerman, 2022; Azabou et al., 2024), but prior works use supervised labels, so do not leverage large amounts of unannotated data. The brain-computer-interface field has studied how to align latent spaces (Pandarinath et al., 2018; Karpowicz et al., 2022; Degenhart et al., 2020; Jude et al., ; Ma et al., 2023) which either still requires creating an alignment matrix to learn across datasets or only provides post-training alignment mechanisms rather than learning across datasets. Other approaches impute missing channels or learn latent spaces robust to missing channels (Talukder et al., 2022; Zhang et al., 2021; Chau et al., 2024), but these are more suited for the occasional missing channel rather than largely varying sensor layouts. We directly learn spatial-level representations using self-supervised learning across datasets to leverage massive amounts of unannotated intracranial data.
Figure 1: Schematic of our approach. The inputs to our model (a) are the combined neural activities from a collection of intracranial electrodes in a given time interval. These are passed to a frozen temporal embedding model (dotted red outline: BrainBERT shown), which produces a set of time-contextual embedding vectors (yellow). The 3D position of each electrode (red) is added to these vectors to produce the model inputs (orange). The PopT produces space-contextual embeddings for each electrode and a [CLS] token (blue), which can be fine-tuned for downstream tasks. During pretraining, the PopT is trained on two objectives simultaneously. In the first, (b) the PopT determines whether two different sets of electrodes (orange vs brown) represent consecutive or non-consecutive times. In the second objective, (c) the PopT must determine whether an input channel has been replaced with activity at a random other time that is inconsistent with the majority of inputs.
@@ 3 Population Transformer Approach
Figure 1 overviews our Population Transformer (PopT) approach. The key ideas are: (1) to learn a generic representation of neural recordings that can handle arbitrary electrode configurations; and (2) to employ a modular system design that uses a transformer architecture to aggregate information from existing per-channel temporal embeddings. To do so, we employ a self-supervised pretraining approach to learn ensemble and channel level representations. Afterwards, one can fine-tune PopT on downstream decoding tasks. In addition to offering strong decoding results, including generalization to new subjects with different electrode configurations than training subjects (see Section 5), the modular system design is computationally lightweight (see Appendix B), can benefit from improved temporal representations, and is more readily interpretable (see Section 6).
@@
Architecture A schematic of our Population Transformer (PopT) approach is shown in Figure 1. We adapt a transformer backbone due to its ability to accommodate variable channel configurations. Consider a given subject with $N$ channels indexed by $C=\{1,...,N_{c}\}$, and an arbitrary subset of channels $S\subseteq C$. Activity from channel $i$ at time $t$ can be denoted by $x^{t}_{i}$. The PopT takes as input an interval of brain activity $X^{t}=\{x^{t}_{i}|i\in S\}$ from a given time $t$ and a special [CLS] token. Per channel, each interval of brain activity is passed through a temporal embedding model $B$, in the figure’s case BrainBERT, to obtain a representation of each channel’s temporal context.
@@
To allow the model to learn a common brain state representation across layouts, each channel’s embedding is augmented with its 3D position, so that the final input to the PopT is $X^{t}_{B}=\{B(x^{t}_{i})+pos(i)+\mathcal{N}(0,\sigma)|x^{t}_{i}\in X^{t}\}$. Spatial location is given by the electrode’s Left, Posterior, and Inferior coordinates for iEEG electrodes (Wideman, 2024), and XYZ positions for EEG electrodes. We add Gaussian fuzzing to each coordinate location to prevent overfitting to a particular set of coordinates. Each coordinate is then encoded using sinusoidal position encoding (Vaswani et al., 2017). Membership in a particular ensemble (see below: ensemble-wise loss) is also encoded. The four encodings are concatenated together to form the position embedding $pos(i)=[e_{\text{left}};e_{\text{post.}};e_{\text{inf}};e_{\text{ensemble}}]$.
@@
The core of PopT consists of a transformer encoder stack (see Appendix A: Architectures). The output of the PopT are spatial-contextual embeddings of the channels $Y=\{y_{i}\}$ as well as an embedding of the CLS token $y_{cls}$. During pretraining, the PopulationTransformer additionally is equipped with a linear layer head for the [CLS] token output and separate linear layer heads for all other individual token outputs. These produce the scalars $\tilde{y}_{cls}$ and $\tilde{y}_{i}$ respectively, which are used in the pretraining objective (see Figure 1b and c).
@@
Self-supervised loss Our loss function has two discriminative components: (1) ensemble-wise — the model determines if activities from two channel ensembles occurred consecutively, requiring an effective brain state representation at the ensemble-level, (2) channel-wise — the model identifies outlier channels that have been swapped with a different timepoint’s activity, requiring sensitivity to surrounding channel context.
@@
A key aspect of our method is the fact that our objective is discriminative, rather than reconstructive, as is often the case in self-supervision (Liu et al., 2021; Wang et al., 2022). We found this to be necessary, because in practice, the temporal embeddings often have low effective dimension (see Wang et al. (2022)), and reconstruction rewards the model for overfitting to “filler” dimensions in the feature vector (see Section 5).
@@
Pretraining In ensemble-wise discrimination (fig. 1b), two different subsets of channels $S_{A},S_{B}\subset C$ are chosen with the condition that they be disjoint $S_{A}\cap S_{B}=\emptyset$. During pretraining, the model receives the activities from these channels at separate times $X_{A}^{t}=\{x^{t}_{i}\mid i\in S_{A}\}$ and $X_{B}^{t^{\prime}}=\{x^{t^{\prime}}_{i}\mid i\in S_{B}\}$. The objective of the task is then to determine whether these states $X_{A}^{t}$ and $X_{B}^{t^{\prime}}$ have occurred consecutively in time or are separated by some further, randomly selected interval. Given the output of the classification head, the loss function $\mathcal{L}_{N}$ is again the binary cross entropy. We also vary the number of input channels during sampling to ensure the model handles ensembles of different sizes.
Additionally, we select disjoint subsets for ensemble-wise discrimination to prevent the model from solving tasks through trivial copying.
@@
In channel-wise discrimination (fig. 1c), the model must determine whether a channel’s activity has been swapped with activity from a random time. Precisely, activity from each channel $i$ is drawn from a time $t_{i}$. All channels are drawn from the same time $t_{i}=T$, and then 10% of the channels are randomly selected to have their activity replaced with activity from the same channel, but taken from a random point in time $t_{i}\neq T$. Then, given the token outputs of PopT, the channel-wise loss function $\mathcal{L}_{C}$ is the binary cross entropy. Then, our complete objective function is $\mathcal{L}=\mathcal{L}_{N}+\mathcal{L}_{C}$.
@@
Fine-tuning During fine-tuning, the [CLS] intermediate representation, $\tilde{y}_{cls}$ of the pretrained PopT is passed through a single layer linear neural network to produce a scalar $\hat{y}_{cls}$; this forms the input to the binary cross entropy loss for our decoding tasks (see Section 4).
@@ 4 Experiment Setup
Data We use two types of neural time-series data: intracranial and scalp electroencepholography (iEEG and EEG). Whereas iEEG probes are surgically implanted within the 3D brain volume, EEG electrodes lie on the scalp. These cover two resolution extremes of neural time-series data modalities.
@@
iEEG: We use the publicly available subject data from Wang et al. (2022). Data was collected from 10 subjects (total 1,688 electrodes, with a mean of 167 electrodes per subject) who watched 26 movies while intracranial probes recorded their brain activity. Of the sessions, 19 are used for pretraining, and 7 of the sessions are held-out for evaluation.
@@
EEG: We use the Temple University Hospital EEG and Abnormal datasets (TUEG and TUAB (Obeid & Picone, 2016)) for pretraining and task data respectively. We remove all task subjects from the pretraining set and follow the data preprocessing practices in Yang et al. (2024); Jiang et al. (2024).
@@
Decoding Tasks We evaluate on 5 different classification tasks: 4 auditory-linguistic tasks used in the evaluation of Wang et al. (2022) and 1 widely evaluated seizure detection task from Obeid & Picone (2016). Of the auditory-linguistic tasks, two of the tasks are audio focused: determining whether a word is spoken with a high or low pitch and determining whether a word is spoken loudly or softly. And two of the tasks have a more linguistic focus: determining whether the beginning of a sentence is occurring or determining whether any speech at all is occurring. The EEG TUAB seizure detection task is a binary classification of pathological or normal EEG recording.
To test scaling with arbitrary ensemble sizes, we select subsets of electrodes based on their individual linear decodability, with the smallest subsets containing the electrodes with highest decodability.
@@
Baselines For controlled baselines, we concatenate the single-channel temporal embeddings and train a linear (Linear) or non-linear (Deep NN) aggregator on the decoding task. These enable us to directly assess how much PopT improves upon existing aggregation approaches (Ghosal & Abbasi-Asl, 2021). These approaches cannot be pretrained across subjects due to the changing meaning and quantity of inputs. To test the effectiveness of pretraining, we also compare against a non-pretrained PopT. Methods compared For the iEEG experiments, we also compare against Brant (Zhang et al., 2024), which is an end-to-end iEEG trained encoder. We take the fully pretrained Brant model, and fine-tune on our iEEG tasks combining channels with linear aggregation. For the EEG experiments, we compare against BIOT (Yang et al., 2024) and LaBraM (Jiang et al., 2024). Both BIOT and LaBraM use discrete channel ID as positional encodings, which only allows them to be used for layouts with discete channel locations such as EEG. They also train end-to-end in contrast to our modular approach.
@@
      Linear Agg. $0.59\pm 0.08$ $0.66\pm 0.08$ $0.70\pm 0.09$ $0.71\pm 0.11$
      Deep NN Agg. $0.58\pm 0.08$ $0.67\pm 0.08$ $0.71\pm 0.10$ $0.72\pm 0.10$
      Non-pretrained PopT $0.53\pm 0.06$ $0.61\pm 0.13$ $0.74\pm 0.10$ $0.70\pm 0.08$
      Pretrained PopT $\mathbf{0.69}\pm\mathbf{0.07}$ $\mathbf{0.84}\pm\mathbf{0.06}$ $\mathbf{0.86}\pm\mathbf{0.05}$ $\mathbf{0.89}\pm\mathbf{0.07}$
@@
      Linear Agg. $0.55\pm 0.02$ $0.66\pm 0.03$ $0.79\pm 0.04$ $0.77\pm 0.05$
      Deep NN Agg. $0.57\pm 0.02$ $0.67\pm 0.03$ $0.78\pm 0.03$ $0.75\pm 0.05$
      Non-pretrained PopT $0.53\pm 0.02$ $0.64\pm 0.02$ $0.79\pm 0.03$ $0.77\pm 0.05$
      Pretrained PopT $\mathbf{0.60}\pm\mathbf{0.02}$ $\mathbf{0.73}\pm\mathbf{0.02}$ $\mathbf{0.86}\pm\mathbf{0.03}$ $\mathbf{0.84}\pm\mathbf{0.06}$
@@
Table 1: Pretraining PopT is critical to downstream decoding performance (iEEG data). We test on a variety of audio-linguistic decoding tasks (see Section 4) with 90 channels as input. The temporal encoder used for aggregation in sections 1 and 2 are denoted in the section header. We also evaluate against an end-to-end pretrained iEEG model in section 3. Shown are the ROC-AUC mean and standard error across subjects. Best per section are bolded.
@@
      Pretrained PopT $\mathbf{0.7976}\pm\mathbf{0.0022}$ $\mathbf{0.8821}\pm\mathbf{0.0016}$
TS2Vec
      Linear Agg. $0.7649\pm 0.0005$ $0.8533\pm 0.0003$
      Deep NN Agg. $0.7853\pm 0.0021$ $0.8721\pm 0.0015$
@@
      Pretrained PopT $\mathbf{0.8063}\pm\mathbf{0.0010}$ $\mathbf{0.8907}\pm\mathbf{0.0019}$

### v4

@@ Abstract
We present a self-supervised framework that learns population-level codes for arbitrary ensembles of neural recordings at scale. We address key challenges in scaling models with neural time-series data, namely, sparse and variable electrode distribution across subjects and datasets. The Population Transformer (PopT) stacks on top of pretrained temporal embeddings and enhances downstream decoding by enabling learned aggregation of multiple spatially-sparse data channels. The pretrained PopT lowers the amount of data required for downstream decoding experiments, while increasing accuracy, even on held-out subjects and tasks. Compared to end-to-end methods, this approach is computationally lightweight, while achieving similar or better decoding performance. We further show how our framework is generalizable to multiple time-series embeddings and neural data modalities. Beyond decoding, we interpret the pretrained and fine-tuned PopT models to show how they can be used to extract neuroscience insights from large amounts of data. We release our code as well as a pretrained PopT to enable off-the-shelf improvements in multi-channel intracranial data decoding and interpretability. Code is available at https://github.com/czlwang/PopulationTransformer.
1California Institute of Technology
{gchau, sabera, ssoedarm, yyue}@caltech.edu
2MIT CSAIL, CBMM
{czw, vsub851, boris, abarbu}@mit.edu
@@ 1 Introduction
Building effective representations of neural data is an important tool in enabling neuroscience research. Recordings from the brain such as intracranial (iEEG) and scalp (EEG) electroencepholography, consist of time series recorded simultaneously from multiple channels. The relationships between these time series are complex and governed by the underlying functional connectivity that exists between brain regions. Our goal is to build an effective model of multi-channel activity. Recently, improvements have been made in modeling time-series (Wang et al., 2022; Talukder et al., 2024; Yue et al., 2022; Ansari et al., 2024). This suggests an approach for learning multi-channel representations via aggregating temporal embeddings. However, this is not a trivial task. For brain recordings, particularly iEEG, one must contend with sparse and variable electrode layouts, which change the semantics of input channels from subject to subject. This forces many Brain Machine Interface (BMI) and neuroscience studies to rely on expensive schemes, in which models are retrained for each new participant, requiring large amounts of data for calibration (Faezi et al., 2021; Herff et al., 2020; Martin et al., 2018; Metzger et al., 2023; Willett et al., 2023). To this end, we propose a self-supervised learning framework, Population Transformer (PopT), which is specifically designed to aggregate single-channel encodings across variable electrode layouts.
Self-supervised pretraining on unannotated data has been shown to be effective for creating generic representations that are useful for many downstream tasks (Bommasani et al., 2022). Prior work has shown how to pretrain subject-specific (Le & Shlizerman, 2022) or channel-specific (Wang et al., 2022) models of iEEG, but such techniques ignore inter-channel relationships or commonalities that might exist across subjects. Recent end-to-end self-supervised learning approaches downsample signals heavily to make training across hundreds of channels feasible (Zhang et al., 2024; Yang et al., 2024; Jiang et al., 2024). This is particularly problematic for high-fidelity iEEG signals, which capture sub-millisecond changes in neural activity. Our approach leverages existing rich temporal embeddings to represent signal, freeing the model to focus on learning effective aggregation.
@@
We propose Population Transformer (PopT), a self-supervised pretraining approach that learns subject-generic representations of arbitrary electrode ensembles. Transformers offer the flexibility to learn aggregate information across channels, but large amounts of data are needed to train the attention weights (Devlin et al., 2019). During pretraining, we train on large amounts of unannotated data and simultaneously optimize both a channel-level and ensemble-level objective. These tasks encourage the model to develop subject-generic representations by modeling (1) individual channels in the context of surrounding channels and (2) channel ensembles and their relations across time.
@@
Our PopT approach is modular, and builds on top of powerful single-channel temporal embeddings, which provides two key advantages. First, by separating single-channel embedding and multi-channel-aggregation into different modules, we make our approach agnostic to the specific type of temporal embedding used, leaving room for future independent improvements along either the temporal or spatial dimension (an approach that has been validated in video modeling (Arnab et al., 2021)). Second, by taking advantage of learned channel embeddings, PopT training is computationally lightweight compared to their end-to-end counterparts (Appendix B) and baseline aggregation approaches (Figure 4), allowing for adoption in lower compute resource environments.
@@
Empirically, we find that our pretrained PopT outperforms commonly used aggregation approaches (Ghosal & Abbasi-Asl, 2021), and is competitive with end-to-end trained methods (Zhang et al., 2024; Yang et al., 2024; You et al., 2019). Moreover, we find that these benefits hold even for subjects not seen during pretraining, indicating its usefulness for new subject decoding. We also show that the pretrained PopT weights themselves reveal interpretable patterns for neuroscientific study. Finally, we demonstrate that our proposed framework is agnostic to the underlying temporal encoder, further allowing it to adapt to other neural recording modalities.
@@ 2 Related Work
Modeling across variable input channels Modeling spatial representations on top of temporal embeddings has been found to be beneficial for decoding (Faezi et al., 2021; Le & Shlizerman, 2022; Azabou et al., 2024), but prior works use supervised labels, so do not leverage large amounts of unannotated data. The brain-computer-interface field has studied how to align latent spaces (Pandarinath et al., 2018; Karpowicz et al., 2022; Degenhart et al., 2020; Jude et al., ; Ma et al., 2023) which either still requires creating an alignment matrix to learn across datasets or only provides post-training alignment mechanisms rather than learning across datasets. Other approaches impute missing channels or learn latent spaces robust to missing channels (Talukder et al., 2022; Zhang et al., 2021; Chau et al., 2024), but these are more suited for the occasional missing channel rather than largely varying sensor layouts. We directly learn spatial-level representations using self-supervised learning across datasets to leverage large amounts of unannotated intracranial data.
@@ 3 Population Transformer Approach
Figure 1: Schematic of our approach. The inputs to our model (a) are the neural activities from a collection of electrodes in a given time interval (bottom). These are passed to a frozen temporal embedding model (dotted red outline: BrainBERT (Wang et al., 2022) shown), which produces a set of time embedding vectors (yellow). The 3D positions of each electrode (red) are summed with these vectors to produce the model inputs (orange, lower). PopT produces space-contextual embeddings (orange, top) for each electrode and a [CLS] token (blue, top), which can be fine-tuned for downstream tasks. In pretraining, PopT learns two objectives simultaneously. In the first, (b) PopT determines whether two different sets of electrodes (orange vs brown) represent consecutive or non-consecutive times. In the second objective, (c) PopT must determine whether an input channel has been replaced with activity at a random other time that is inconsistent with the majority of inputs.
@@
Figure 1 overviews our Population Transformer (PopT) approach. The key ideas are: (1) to learn a generic representation of neural recordings that can handle arbitrary electrode ensembles; and (2) to employ a modular system design that uses a transformer architecture to aggregate information from existing per-channel temporal embeddings. To do so, we employ a self-supervised pretraining approach to learn ensemble and channel level representations. Afterwards, one can fine-tune PopT on downstream decoding tasks. In addition to offering strong decoding results, including generalization to new subjects with different electrode configurations than training subjects (see Section 5), the modular system design is computationally lightweight (see Appendix B), can benefit from improved temporal representations, and is more readily interpretable (see Section 6).
@@
Architecture A schematic of our Population Transformer (PopT) approach is shown in Figure 1. We adopt a transformer backbone due to its ability to accommodate variable channel configurations. Consider a given subject with $N$ channels indexed by $C=\{1,...,N_{c}\}$, and an arbitrary subset of channels $S\subseteq C$. Let $x^{t}_{i}\in\mathbb{R}^{T}$ denote a time window of activity from channel $i$ that begins at time $t$, where $T$ is the number of time samples in the interval. The PopT takes as input a collection of such channels activities, $X^{t}=\{x^{t}_{i}|i\in S\}$, as well as a special [CLS] token. Per channel, each interval of brain activity is passed through a temporal embedding model $B$, in the figure’s case BrainBERT (Wang et al., 2022), to obtain a representation of each channel’s temporal context, $B(x^{t}_{i})\in\mathbb{R}^{d}$, where $d$ is the embedding dimension. For BrainBERT, the first step of pre-processing involves obtaining the STFT for the signal, but preprocessing will differ depending on the embedding model used.
@@
To allow the model to learn a common brain state representation across layouts, each channel’s embedding is summed with its 3D position, so that the final processed input to the PopT is ${X^{t}_{B}=\{B(x^{t}_{i})+pos(i)|x^{t}_{i}\in X^{t}\}}$. The PopT receives this as an $S\times d$ matrix. Spatial location is given by the electrode’s Left, Posterior, and Inferior coordinates for iEEG electrodes (Wideman, 2024), and XYZ positions for EEG electrodes. Membership in a particular ensemble (see below: ensemble-wise loss) is also encoded. The four encodings are concatenated together to form the position embedding $pos(i)=[e_{\text{left}};e_{\text{post.}};e_{\text{inf}};e_{\text{ensemble}}]$, where $e$ is given using a sinusoidal position encoding that represents a scalar coordinate as a unique combination of sines (Vaswani et al., 2017).
@@
The core of PopT consists of a transformer encoder stack (see Appendix A: Architectures). The output of the PopT are spatial-contextual embeddings of the channels $Y=\{y_{i}\}$ and an embedding of the CLS token $y_{cls}$. During pretraining, the PopT additionally is equipped with a linear head for the [CLS] token output and separate linear heads for all other individual token outputs. These produce the scalars $\tilde{y}_{cls}$ and $\tilde{y}_{i}$ respectively, which are used in the pretraining objective (Figure 1b and c).
@@
Self-supervised loss Our loss has two discriminative components: (1) ensemble-wise — the model determines if activity from two ensembles occurred consecutively, requiring an effective brain state representation at the ensemble-level, (2) channel-wise — the model identifies outlier channels swapped with a different timepoint’s activity, requiring sensitivity to surrounding channel context.
@@
A key aspect of our method is the fact that our objective is discriminative, rather than reconstructive, as is often the case in self-supervision (Liu et al., 2021; Wang et al., 2022). In practice, the temporal embeddings often have low effective dimension (see Wang et al. (2022)), and reconstruction rewards the model for overfitting to “filler” dimensions in the feature vector (Section 5).
@@
Pretraining In ensemble-wise discrimination (fig. 1b), two different subsets of channels $S_{A},S_{B}\subset C$ are chosen with the condition that they be disjoint $S_{A}\cap S_{B}=\emptyset$. During pretraining, the model receives the activities from these channels at separate times $X_{A}^{t}=\{x^{t}_{i}\mid i\in S_{A}\}$ and $X_{B}^{t^{\prime}}=\{x^{t^{\prime}}_{i}\mid i\in S_{B}\}$. The objective of the task is then to determine whether these states $X_{A}^{t}$ and $X_{B}^{t^{\prime}}$ have occurred consecutively in time $(|t-t^{\prime}|=500ms)$ or are separated by some further, randomly selected interval. Given the output of the classification head, the loss function $\mathcal{L}_{N}$ is the binary cross-entropy (BCE). We select disjoint subsets for ensemble-wise discrimination to prevent the model from solving tasks through trivial copying. We also randomly vary $|S|$ during sampling to ensure the model handles ensembles of different sizes.
@@
In channel-wise discrimination (fig. 1c), the model must determine whether a channel’s activity has been swapped with activity from a random time. Precisely, activity from each channel $i$ is drawn from a time $t_{i}=t$ for all channels. Then, 10% of the channels are randomly selected to have their activity replaced with activity from a randomly selected channel at a random point in time $t_{i}\neq t$. For each of the token outputs of PopT, the channel-wise loss function $\mathcal{L}_{C}$ is the BCE. Our complete objective function is $\mathcal{L}=\mathcal{L}_{N}+\mathcal{L}_{C}$. A detailed formulation of the objective is given in Appendix A.
@@
Fine-tuning In fine-tuning, given the [CLS] token, which is a $d$-dimensional vector, the PopT produces the intermediate representation, $\tilde{y}_{cls}\in\mathbb{R}^{d}$, which is passed through a single layer linear to produce a scalar prediction $\hat{y}_{cls}\in\mathbb{R}$. BCE loss is used for our binary decoding tasks (Section 4).
@@ 4 Experiment Setup
Data We use two types of neural time-series data: intracranial and scalp electroencepholography (iEEG and EEG). iEEG probes are surgically implanted within the 3D brain volume and record local electric signals from the brain at very high temporal resolution and spatial precision. EEG electrodes lie on the scalp, and record electric signals that are smeared by the skull, which results in low temporal and spatial resolution. EEG montages typically tile the whole scalp, while iEEG electrodes are often only inserted in a comparatively smaller number of locations. These cover two resolution extremes of neural time-series data modalities.
@@
iEEG: We use the publicly available subject data from Wang et al. (2024). Data was collected from 10 subjects (total 1,688 electrodes, with a mean of 167 electrodes per subject) who watched 26 movies (19 for pretraining, 7 for downstream decoding) while intracranial probes recorded their brain activity. To test decoding with arbitrary ensemble sizes, we select subsets of electrodes based on their individual linear task decodability, with the smallest subsets containing the electrodes with highest decodability. We follow the trialization and data preprocessing practices used in Wang et al. (2022).
@@
EEG: We use the Temple University Hospital EEG and Abnormal datasets, TUEG and TUAB (Obeid & Picone, 2016), for pretraining and task data respectively. We remove all task subjects from the pretraining set and follow the data preprocessing practices in Yang et al. (2024); Jiang et al. (2024).
@@
Decoding Tasks We evaluate on 5 different classification tasks: 4 auditory-linguistic tasks used in the evaluation of Wang et al. (2022) and 1 widely evaluated abnormal EEG detection task from Obeid & Picone (2016). Of the auditory-linguistic tasks, two of the tasks are audio focused: determining whether a word is spoken with a high or low pitch and determining whether a word is spoken loudly or softly. And two of the tasks have a more linguistic focus: determining whether the beginning of a sentence is occurring or determining whether any speech at all is occurring. The TUAB abnormal EEG detection task is a binary classification of pathological or normal EEG recording.
@@
Baselines For controlled baselines, we concatenate the single-channel temporal embeddings and train a linear (Linear) or non-linear (Deep NN) aggregator on the decoding task. These enable us to directly assess how much PopT improves upon existing aggregation approaches (Ghosal & Abbasi-Asl, 2021). These approaches cannot be pretrained across subjects due to the changing meaning and quantity of inputs. To test the effectiveness of pretraining, we also compare against a non-pretrained PopT.
Methods compared For the iEEG experiments, we also compare against Brant (Zhang et al., 2024), which is an end-to-end iEEG encoder. We take the fully pretrained Brant model, and fine-tune on our iEEG tasks combining channels with linear aggregation. For the EEG experiments, we compare against reported BIOT (Yang et al., 2024) and LaBraM (Jiang et al., 2024) results. They also train both temporal and spatial encoders together in contrast to our modular approach.
Temporal encoders To test the generalizability of our approach, we train with a variety of temporal encoders: BrainBERT (Wang et al., 2022), which is designed for iEEG data, TOTEM (Talukder et al., 2024) which learns a tokenization of the input, Chronos (Ansari et al., 2024) which is a large general time-series encoder, and TS2Vec (Yue et al., 2022) which has a hierarchical convolutional architecture. Hidden dimensions of these encoders vary from 64 to 768. More details are in Appendix A.
@@
      Linear Agg. $0.59\pm 0.03$ $0.66\pm 0.03$ $0.70\pm 0.04$ $0.71\pm 0.04$
      Deep NN Agg. $0.56\pm 0.03$ $0.64\pm 0.04$ $0.71\pm 0.03$ $0.70\pm 0.04$
      Non-pretrained PopT $0.53\pm 0.02$ $0.61\pm 0.05$ $0.74\pm 0.04$ $0.70\pm 0.03$
      Pretrained PopT $\mathbf{0.74}\pm\mathbf{0.03}^{*}$ $\mathbf{0.87}\pm\mathbf{0.03}^{*}$ $\mathbf{0.90}\pm\mathbf{0.01}^{*}$ $\mathbf{0.93}\pm\mathbf{0.02}^{*}$
@@
      Linear Agg. $0.57\pm 0.02$ $0.67\pm 0.02$ $0.80\pm 0.03$ $0.77\pm 0.05$
      Deep NN Agg. $0.58\pm 0.02$ $0.70\pm 0.02$ $0.80\pm 0.03$ $0.77\pm 0.05$
      Non-pretrained PopT $0.53\pm 0.01$ $0.62\pm 0.02$ $0.80\pm 0.03$ $0.77\pm 0.05$
      Pretrained PopT $\mathbf{0.64}\pm\mathbf{0.03}$ $\mathbf{0.79}\pm\mathbf{0.02}^{*}$ $\mathbf{0.90}\pm\mathbf{0.02}^{*}$ $\mathbf{0.88}\pm\mathbf{0.05}^{*}$
@@
Table 1: Pretraining PopT is critical to downstream decoding performance (iEEG data). We test on a variety of audio-linguistic decoding tasks (see Section 4) with 90 channels as input. The temporal encoder used for aggregation in sections 1 and 2 are denoted in the section header. We also evaluate against an end-to-end pretrained iEEG model in section 3. Shown are the ROC-AUC mean and standard error across subjects. Best per section are bolded. Asterisks indicate that the bolded model is significantly better than the second-place model ($p<0.05$, Wilcoxon rank-sum).
@@
      Pretrained PopT $\mathbf{0.7963}\pm\mathbf{0.0021}^{*}$ $\mathbf{0.8822}\pm\mathbf{0.0013}^{*}$
TS2Vec:
      Linear Agg. $0.7554\pm 0.0005$ $0.8549\pm 0.0011$
      Deep NN Agg. $0.7835\pm 0.0039$ $0.8723\pm 0.0029$
@@
      Pretrained PopT $\mathbf{0.8060}\pm\mathbf{0.0025}^{*}$ $\mathbf{0.8883}\pm\mathbf{0.0008}^{*}$